# Hyperparameter Optimization in Machine Learning

### v1

@@ 5.2.3 Hyperband
The minimum budget $b_{min}$ allocated to each configuration is an important meta-parameter of SH. If $b_{min}$ is too small, we may filter out good configurations based on unreliably low fidelities, while for too large $b_{min}$, our edge over standard random search may diminish. Hyperband [li2017hyperband] deals with this problem by running different instances $l=0,\dots,K$ of SH in sequence. The $l$-th instance (or bracket) has rung levels $\mathcal{R}_{l}=\{b_{min}\eta^{k}\;|\;k=l,\dots,K\}$, so $l=0$ corresponds to SH with $K+1$ rung levels, while $l=K$ is random search with a single level $b_{max}$. Each bracket has its own number of initial configurations $n_{l}$, which is chosen such that the total budget spent in each bracket is roughly the same. The risk of choosing too small a value for $b_{min}$ is mitigated at modest extra expense.
What about parallel computation? Being based on SH, Hyperband inherits the issue of idle workers noted above for synchronous scheduling. However, since brackets are independent of each other, we can work on several of them in parallel, which (at the expense of some implementation complexity) eliminates idle time for asynchronous scheduling.
@@ 5.2.4 Asynchronous successive halving (ASHA)
As noted above, neither successive halving nor Hyperband fare well in the presence of parallel computation. Moreover, even if a configuration is clearly superior to others in all fidelities, it takes a long time until it is promoted to the maximum level $b_{max}$. In fact, before any configuration can be promoted towards the next level, all slots in the current rung have to be filled. These shortcomings are addressed in asynchronous successive halving (ASHA) [Li:19]. In this algorithm, rungs do not have a predefined size, but only grow over time. Also, promotion decisions are done less conservatively. For example, assume that $\eta=3$. Scanning rungs from highest to lowest, any not-yet-promoted configuration whose metric value is smaller than $2/3$ of other entries, is promoted to the next level. If no configuration can be promoted, a new one is sampled for the lowest level. In other words, configurations are promoted as early as possible: once a rung has 3 entries, one of them moves to the next level. In ASHA, any available compute resource can be used instantly, and workers are never idle. Moreover, any superior configuration is rapidly promoted towards $b_{max}$, without having to wait at a number of synchronization points. In practice, ASHA tends to outperform synchronous SH or Hyperband significantly [Klein:20].
@@ 5.3 Model-based methods
So far, the multi-fidelity methods we looked at sample hyperparameter configurations purely at random. As we have seen in Chapter 4, more sophisticated methods that use a probabilistic model to sample configurations usually require less function evaluations to approach the global optimum than random-search-based approaches. In this section, we will discuss multi-fidelity strategies that use one or multiple probabilistic models to guide the search.
@@ 5.3.1 Fabolas
Fabolas [klein2017fast] models the performance $f_{b}(\lambda)$ of a hyperparameter configuration $\lambda$ across different subsets $b\in[b_{min},b_{max}]$ of the training data. Compared to classical Bayesian optimization (see Chapter 4), in this case, we cannot simply optimize standard acquisition functions, such as expected improvement, because they tend to over-exploit in this setting and mostly evaluate configurations on the highest budget.
@@
Instead, Fabolas is based on the entropy search method by Hennig2012 which models the probability
$p_{\min}(\lambda\mid\mathcal{D})=p(\lambda\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f(\lambda))$
@@
of $\lambda$ being the global optimum. Now, the idea of entropy search is to select in each iteration the configuration that provides the most information about $p_{min}$, which we quantify by computing the KL-Divergence $KL\big{[}p_{min}(\lambda\mid\mathcal{D})||\mathcal{U}(\lambda)\big{]}$ between $p_{min}$ and a uniform distribution $\mathcal{U(\lambda})$ which serves as a prior positing that each point $\lambda\in\Lambda$ is equally likely to be the optimum. In other words, entropy search selects in each iteration the candidate point $\lambda$ that mostly moves $p_{min}$ away from a Uniform to a highly peaked distribution around the global optimum.
@@
For Fabolas, we model $p_{min}$ only at the highest $b=b_{max}$, since we aim to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$ Now, to balance between information gain and the cost of evaluating $\lambda$, Fabolas maintains a second Gaussian process $p(c\mid\lambda,b)$ to model the wall-clock time $c(\lambda,b)$ of evaluating $f_{b}(\lambda)$. In each iteration, Fabolas optimizes the following acquisition function:
$a(\lambda,b)=\frac{KL\big{[}p_{min}(\lambda\mid\mathcal{D})||\mathcal{U}(\lambda)\big{]}}{\mu_{c}(\lambda,b)}$
@@
to select both the configuration $\lambda$ and budget $b$ that provides the most information about $p_{min}$ while being not too expensive to evaluate, where $\mu_{c}(\lambda,b)=\mathbb{E}_{p(c\mid\lambda,b)}[c(\lambda,b)]$ is the mean predicted cost of evaluating $\lambda$ using a budget $b$.
5.3.2 BO-HB
To combine the sample efficiency of Bayesian optimization with the strong anytime performance of Hyperband, falkner2018bohb fits an independent TPE model (see Section 4.3) on each rung level $b_{i}\in\mathcal{R}$ to sample new configurations. More specifically, once the proposed method BO-HB observes $N_{min}$ datapoints on the rung level $b_{i}$, it trains two kernel density estimator as shown in Eq. (4.3) on all data points of this rung level. New configurations are now sampled according to density ratio $l(\lambda)/g(\lambda)$ instead of uniformly at random. This continues until we observed $N_{min}$ on the highest rung level $b_{max}$ and we keep sampling from the model on this rung level. While the original implementation of TPE by bergstra2011algorithms used uni-variant kernel density estimator, falkner2018bohb proposed to use multi-variant kernel density estimator to capture the interaction between hyperparameters. To fit such a model, we need to have at least $N_{min}=m+1$ datapoints where $m$ is the dimensionality of the search space.
@@ 5.3.2 BO-HB
Figure 5.3 shows a comparison by [falkner2018bohb] of BO-HB to other approaches for optimizing the hyperparameters of a neural network. Hyperband outperforms in the beginning, as it is able to return configurations much earlier than non-multi-fidelity methods such as random search or TPE. However, compared to model-based approaches such as TPE, Hyperband eventually converges more and more to random search, as it samples configurations uniformly at random and, in later stages, requires evaluating them on the larger budgets to improve. BO-HB, on the other hand, maintains the early speed-up of Hyperband but achieves a similar strong performance than TPE in the later stages of the optimization process.
@@ 6.1 Evolutionary algorithms
The terminology of EA techniques is borrowed from biology: the set of points is called population and candidates are called individuals, chromosomes or genotypes. Iterations are called generations and the objective function is called fitness. Typically at the end of each generation the fitness of all individuals is observed, making EA inherently parallel algorithms. Traditional implementations typically also require synchronization. After observing the fitness of the population a series of transformations is applied to the individuals. Common types of transformations, which collectively implement explore/exploit strategies in this context, include:
@@
  • Recombination or crossover, global transformations that consist in generating new individuals combining slices of two or more parents;
@@
Other strategies like elitism (keep the best individuals unchanged) and niching (maintain a diverse population) are often used in conjunction.
Various implementations of ES/GA have been proposed in HPO e.g. for tuning hyperparameters of support vector machines [friedrichs2005evolutionary], for neural architecture search [real2019regularized, miikkulainen2024evolving] (see Section 8.3 for more details), parameters of games [kunanusont2017ntuple] and even entire machine learning algorithms [real2020automl]. In the next subsections, we review two particular implementations that have received notable attention.
6.1.1 Covariance matrix adaptation evolution strategy (CMA-ES)
CMA-ES [hansen1996adapting, hansen2016cma, loshchilov2016cma] maintains a multivariate Gaussian distribution from which it samples individuals for the new generations. That is, the generation at iteration $k+1$ is drawn according to
$\lambda_{k+1}^{j}\sim\mu_{k}+\sigma_{k}\mathcal{N}(0,\Sigma_{k})\quad\text{for}j\in[N]$
where $N\in\mathbb{N}^{+}$ is the population size, $\mu_{k}$ and $\Sigma_{k}$ are the distribution parameters and $\sigma_{k}>0$ is an overall standard deviation, akin to a step size.
The parameters of the distribution are evolved at each iteration following a selection and recombination principle. Suppose $\{\lambda_{k}^{j}\}_{j=1}^{N}$ are sorted in increasing objective value, i.e. $f(\lambda_{k}^{1})\leq f(\lambda_{k}^{2})\leq\dots\leq f(\lambda_{k}^{N})$. Then the mean is updated as a convex combination of the best $1\leq M\leq N$ points:
@@ 6.1.1 Covariance matrix adaptation evolution strategy (CMA-ES)
where the $w_{i}$ are the recombination weights. In practice the update of Eq. (6.1) could be unreliable especially for small populations and hansen2003reducing proposes to maintain a moving average for the covariance matrix. Current implementation of CMA-ES incorporate other strategies to improve robustness and efficiency of the algorithms like adaptation of the step-size $\sigma_{k}$ and generalizations of Eq. (6.1) [hansen2019pycma].
6.1.2 Population-based training (PBT)
@@
jaderberg2017population propose an asynchronous method that evolves a population of models with different hyperparameter configurations as they are trained. The procedure, devised in the context of learning deep neural networks in challenging scenarios like image generation or reinforcement learning, consists in interleaving neural network training (by standard gradient descent) with applications of evolutionary transformations to both the parameter of the models and the hyperparameter of the learning algorithm. PBT is an instance of online HPO which construct an hyperparameter schedule rather than finding a single configuration. We will discuss in more generality hyperparameter schedules and online HPO in Section 8.2.
Let $A=A_{T}\circ\dots\circ A_{1}$ be a decomposition of the base learning algorithm [^1] Such decomposition is exact when each $A_{t}$ is applied with the same, fixed, hyperparameter configuration.. Each $A_{t}$ could be, for instance, an epoch of gradient-based training of a neural network. Let $N$ be a population size and $\{(\theta_{1}^{j},\lambda_{1}^{j})\}_{k\in[N]}$ be a set of pairs of model parameters and algorithm hyperparameters — the individuals in PBT — initialized from suitable distributions. At each generation, individuals are updated (asynchronously) as
$\theta_{k}^{t+1}=A(\theta_{k}^{t},\lambda^{t}_{k})\quad\text{for }k\in[N].$
After this, some evolutionary transformations are applied to the pairs of parameters and hyperparameters using as fitness the evaluation of the HPO objective $\mathscr{M}$ on the hypotheses provided by the $\theta_{k}^{t+1}$’s. jaderberg2017population propose two type of operations:
@@
  • “exploit,” a selection strategy that involves copying over parameters and hyperparameters from other better performing individuals, for instance, by replacing the worst-performing 20% with randomly drawn individuals from the best-performing 20% (truncation selection);
  • “explore,” a mutation tranformation that either perturbs the hyperparameter configuration or resamples it from the initial distribution. In both cases, the parameters are left unvaried.
@@
PBT further uses elitism. tao2020learning combine PBT with gradient-based optimization of the hyperparameters, which we will discuss in the next chapter.

### v2

@@ 5.2.2 Hyperband
The minimum budget $b_{min}$ allocated to each configuration is an important meta-parameter of SH. If $b_{min}$ is too small, we may filter out good configurations based on unreliably low fidelities, while if $b_{min}$ is too large, our advantage over standard random search may diminish. Hyperband [li2017hyperband] addresses this issue by running different instances, $l=0,\dots,K$, of SH in sequence. The $l$-th instance (or bracket) has rung levels
$\mathcal{R}_{l}=\{b_{min}\eta^{k}\mid k\in\{l,\dots,K\}\},$
so $l=0$ corresponds to SH with $K+1$ rung levels, while $l=K$ corresponds to random search with a single level $b_{max}$. Each bracket has its own number of initial configurations $n_{l}$, which is chosen such that the total budget spent in each bracket is roughly the same. The risk of choosing too small a value for $b_{min}$ is mitigated at a modest extra expense.
5.2.3 Synchronous successive halving
How is parallel computation exploited by SH? Recall the notation of synchronous versus asynchronous parallelization from Section 2.4, and assume that $W$ workers are available. For synchronous scheduling, we can start $\min(W,M)$ jobs in parallel, where $M$ is the number of free slots remaining in the currently processed rung. Here, $W-M$ workers remain idle, which is particularly wasteful for the higher rungs. For example, if $n=\eta^{K}$, the highest rung has only a single slot, so $W-1$ workers will be idle for the maximum budget $b_{max}$. The same issue arises for asynchronous scheduling. A remedy will be discussed in Section 5.2.4.
Hyperband inherits the idle worker issue noted above for synchronous scheduling. However, since brackets are independent of each other, we can work on several of them in parallel, which (at the expense of some implementation complexity) eliminates idle time for asynchronous scheduling.
@@ 5.2.4 Asynchronous successive halving (ASHA)
As discussed above, neither SH nor Hyperband can fully exploit parallel computation because they require synchronization among workers evaluating hyperparameter configurations for a specific run.
This also means that if a configuration is clearly superior to all others on a rung $r_{i}\in\mathcal{R}$, it may take a long time before it is promoted to the next rung level $r_{i+1}$. In fact, all configurations on the current rung $r_{i}$ must be evaluated before any promotion decisions can be made.
For example, consider Figure 5.3. In a distributed setting, we refer to the evaluation of a hyperparameter configuration as a trial, and in this scenario, we have two workers evaluating four different trials on Rung 0. If Trial-2 takes significantly longer than the others, worker 1 remains idle until all trials on Rung 0 are completed, delaying progress to Rung 1.
Figure 5.3: Scheduling behavior of Synchronous SH with two workers. A darker shade of blue indicates better-performing trials. Some trials may take significantly longer than others, resulting in idle workers until all trials in the current rung are completed.
These shortcomings are addressed by Asynchronous Successive Halving (ASHA) [Li:19]. In this algorithm, rungs do not have a predefined size but grow dynamically over time. Additionally, promotion decisions are made more aggressively.
For example, assume that $\eta=2$; a configuration is immediately promoted to the next rung if it ranks in the top $50\%$ of the current rung. In other words, configurations are promoted as early as possible: once a rung has two entries, one of them advances to the next level. Figure 5.4 illustrates this process: as soon as Trial-1 is completed, evaluation of Trial-0 continues on Rung 1—assuming Trial-0 performs better than Trial-1 — rather than evaluating Trial-3 on Rung 0.
After evaluating Trial-3, no configurations can be promoted because, at this point, only two configurations have been evaluated on Rung 0, and one on Rung 1. As a result, the evaluation proceeds with a new trial at the lowest level.
Figure 5.4: Scheduling behavior of ASHA. As in Figure 5.3, a darker shade of blue indicates better-performing trials. ASHA eliminates idle time by immediately promoting a trial to the next rung as soon as it meets the promotion criteria, rather than waiting for all trials in the current rung to be completed.
In ASHA, all available computational resources are utilized immediately, ensuring that workers are never idle. Additionally, superior configurations are quickly promoted toward $b_{max}$ without the need to wait for synchronization points. In practice, ASHA often significantly outperforms synchronous SH or Hyperband [Klein:20].
@@ 5.3 Model-based methods
So far, the multi-fidelity methods we have examined sample hyperparameter configurations purely at random. As we saw in Chapter 4, more sophisticated methods that use a probabilistic model to sample configurations typically require fewer function evaluations to approach the global optimum compared to random search-based approaches. In this section, we will discuss multi-fidelity strategies that leverage one or multiple probabilistic models to guide the search.
@@ 5.3.1 Fabolas
Fabolas [klein2017fast] models the performance $f_{b}(\lambda)$ of a hyperparameter configuration $\lambda$ across different subsets $b\in[b_{min},b_{max}]$ of the training data.
Compared to classical Bayesian optimization (see Chapter 4), we cannot simply optimize standard acquisition functions as described in Section 4.2.3. For example, in the case of expected improvement, the highest improvements are expected at $b=b_{max}$. As a result, this approach would lead to over-exploitation, primarily evaluating configurations at the highest budget.
@@
Instead, Fabolas is based on the entropy search method by Hennig2012, which models the probability
$p_{\min}(\lambda^{*}\mid\mathcal{C})=\mathbb{P}(\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f(\lambda))$
@@
of $\lambda$ being the global optimum, where $\mathcal{C}$ are the collected observations (see Section 4.2). The key idea of entropy search is to select, in each iteration, the configuration that provides the most information about $p_{\min}$, which we quantify by computing the KL divergence $KL\big{[}p_{\min}(\lambda\mid\mathcal{C})||\mathcal{U}(\lambda)\big{]}$ between $p_{\min}$ and a uniform distribution $\mathcal{U}(\lambda)$, which serves as a prior assuming that each point $\lambda\in\Lambda$ is equally likely to be the optimum. In other words, entropy search selects, in each iteration, the candidate point $\lambda$ that shifts $p_{\min}$ from a uniform distribution to a highly peaked distribution around the global optimum.
@@
For Fabolas, we model $p_{\min}$ only at the highest budget $b=b_{max}$, since we aim to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$. To balance information gain with the cost of evaluating $\lambda$, Fabolas maintains a second Gaussian process $p(c\mid\lambda,b)$ to model the wall-clock time $c(\lambda,b)$ required to evaluate $f_{b}(\lambda)$. In each iteration, Fabolas optimizes the following acquisition function:
$\alpha(\lambda,b)=\frac{KL\big{[}p_{\min}(\lambda\mid\mathcal{C})||\mathcal{U}(\lambda)\big{]}}{\mu_{c}(\lambda,b)}$
@@
to select both the configuration $\lambda$ and budget $b$ that provide the most information about $p_{\min}$ while remaining cost-effective to evaluate, where $\mu_{c}(\lambda,b)=\mathbb{E}_{p(c\mid\lambda,b)}[c(\lambda,b)]$ is the expected cost of evaluating $\lambda$ using budget $b$.
5.3.2 BO-HB: Bayesian Optimization Hyperband
To combine the sample efficiency of Bayesian optimization with the strong anytime performance of Hyperband, falkner2018bohb fit an independent TPE model (see Section 4.2.5) on each rung level $b_{i}\in\mathcal{R}$ to sample new configurations.
More specifically, once the proposed method BO-HB observes $N_{\min}$ data points at the rung level $b_{i}$, it trains two kernel density estimators—$l(\lambda)$ and $g(\lambda)$—to model the probability of a configuration $\lambda$ achieving a validation error lower or higher than a certain threshold, as defined in Eq. 4.16. However, the density estimators are only trained on all data points at the given rung level.
New configurations are then sampled according to the density ratio $l(\lambda)/g(\lambda)$ instead of being chosen uniformly at random. This process continues until $N_{\min}$ observations have been collected at the highest rung level $b_{max}$, after which new configurations are sampled exclusively from the model trained at this level.
While the original implementation of TPE by bergstra2011algorithms used a univariate kernel density estimator, falkner2018bohb proposed using a multivariate kernel density estimator to better capture interactions between hyperparameters. To fit such a model, at least $N_{\min}=m+1$ data points are required, where $m$ is the dimensionality of the search space.
@@ 5.3.2 BO-HB: Bayesian Optimization Hyperband
Figure 5.5 presents a comparison by falkner2018bohb of BO-HB with other approaches for optimizing the hyperparameters of a neural network. Hyperband initially outperforms other methods since it returns configurations much earlier than non-multi-fidelity methods such as random search or TPE. However, compared to model-based approaches like TPE, Hyperband eventually converges toward random search, as it samples configurations uniformly at random and, in later stages, requires evaluating them on larger budgets to improve performance. BO-HB, on the other hand, retains the early speed-up advantage of Hyperband while achieving a performance level similar to TPE in the later stages of the optimization process.
@@ 6.1 Evolutionary algorithms
The terminology of EA techniques is borrowed from biology: the set of points is called population and configurations are called individuals, chromosomes or genotypes. Iterations are called generations and the objective function is called fitness. Typically at the end of each generation the fitness of all individuals is observed, making EA inherently parallel algorithms – traditional implementations usually requiring synchronization. After observing the fitness of the population, a series of transformations is applied to the individuals. Common types of transformations, which collectively implement explore/exploit strategies in this context, include:
@@
  • Recombination or crossover, global transformations that consist in generating new individuals combining slices (e.g. subsets of hyperparameter coordinates) of two or more parents;
@@
Other strategies like elitism (keep the best individuals unchanged) and niching (maintain a diverse population) are often used in conjunction to the explore/exploit strategies outlined above.
We described this class of algorithm by the procedure outlined in Population-Based-HPO in Algorithm 3. W can think about individuals and generations as draws from $N$ distributions $p_{k}^{j}$, where $N$ is the population size. These distributions implement the strategies described above, and evolve at each generation using global information from the current generation. For instance, elitism can be represented by setting $p_{k}^{j}=\delta_{\lambda_{k}^{j}}$ for the best performing configurations, where $\delta$ is a Dirac delta distribution. In contrast to random or quasi-random search methods covered in Chapter 3, population-based methods keep in general $N$ distinct distributions (rather than a global one) that are updated iteratively (rather than fixed).
Algorithm 3 Population-Based-HPO
1:Response function $f$
2:Population size $N$ and number of generations $T$
3:Initial distributions $\{p^{j}_{0}\}_{j\in[N]}$
4:for $k\leftarrow 1$ to $T$ do
5:     for $j\leftarrow 1$ to $N$ do $\triangleright$ Typically executed in parallel
6:         $\lambda_{k}^{j}\sim p^{j}_{k-1}$triangleright Draw an individual
7:         Evaluate $f(\lambda_{k}^{j})$triangleright Evaluate its fitness (if needed)
8:     end for
9:     Update $p^{j}_{k}$ from $p^{j}_{k-1}$ using $\{(\lambda^{i}_{k},f(\lambda^{i}_{k}))\}_{i\in[N]}$triangleright “Evolve” distributions
10:end for
Various implementations of ES/GA have been proposed in HPO e.g. for tuning hyperparameters of support vector machines [friedrichs2005evolutionary], for neural architecture search [real2019regularized, miikkulainen2024evolving] (see Section 8.3 for a discussion on neural architecture search), parameters of games [kunanusont2017ntuple] and even entire machine learning algorithms [real2020automl]. In the next sections, we review three particular implementations that have received notable attention.
6.2 Covariance matrix adaptation evolution strategy (CMA-ES)
CMA-ES [hansen1996adapting, hansen2016cma, loshchilov2016cma] maintains a global multivariate Gaussian distribution from which it samples individuals for the new generations. That is, the generation at iteration $k+1$ is drawn according to
$\lambda_{k+1}^{j}\sim p_{k}^{j}=\mu_{k}+\sigma_{k}\mathcal{N}(0,\Sigma_{k})\quad\text{for }j\in[N]$
where $N\in\mathbb{N}^{+}$ is the population size, $\mu_{k}$ and $\Sigma_{k}$ are the distribution parameters and $\sigma_{k}>0$ is an overall standard deviation, akin to a step size. Referring to Algorithm 3, this means in CMA-ES $p^{j}_{k}=\mathcal{N}(\mu_{k},\sigma_{k}^{2}\Sigma_{k})$ for all $j\in[N]$.
The parameters of the distribution are evolved at each iteration following a selection and recombination principle. Suppose $\{\lambda_{k}^{j}\}_{j=1}^{N}$ are sorted in increasing objective value, i.e. $f(\lambda_{k}^{1})\leq f(\lambda_{k}^{2})\leq\dots\leq f(\lambda_{k}^{N})$.  [^1] We recall we seek to minimize $f$, hence the first points are the best points. Then the mean is updated as a convex combination of the best $1\leq M\leq N$ points:
@@ 6.2 Covariance matrix adaptation evolution strategy (CMA-ES)
where the $w_{i}$ are the recombination weights. In practice the update of Eq. (6.1) could be unreliable especially for small populations and hansen2003reducing proposes to maintain a moving average for the covariance matrix. Current implementation of CMA-ES incorporate other strategies to improve robustness and efficiency of the algorithms like adaptation of the step-size $\sigma_{k}$ and generalizations of Eq. (6.1); see [hansen2019pycma].
6.3 Population-based training (PBT)
Figure 6.1: Evolution of models trained with PBT for Atari game-playing policies. Red dots represent initial points in the parameter and hyperparameter spaces. Paths represents progressions via parameter training. Branching points represent applications of the exploit and explore operations, namely, a copy of the parameters and hyperparamters of better performing individuals onto worse performing one and a subsequent perturbation of their hyperparameter configurations. The performance score is color-coded, with light blue color representing better performing models. Reproduced from [jaderberg2017population].
@@ 6.3 Population-based training (PBT)
jaderberg2017population propose a parallel method that maintains a population of model and hyperparameter pairs that are evolved as the models are trained. The procedure, devised in the context of learning deep neural networks in challenging scenarios like image generation or reinforcement learning, consists in interleaving neural network training (by gradient descent) with applications of evolutionary transformations to both the parameter of the models and the hyperparameter of the learning algorithm. PBT is an online hyperparamter optimization algorithm, as hyperparameters are updated alongside model parameters. In other words, in contrast to many techniques we discussed so far, PBT yields an hyperparameter sequence (or schedule) rather than a single hyperparamter configuration. 222 We discuss in more generality hyperparameter schedules and online HPO in Section 8.2.
Suppose one can “decompose” the learning algorithm of interest into a series of stages: $A=A_{T}\circ\dots\circ A_{1}$, where each map $A_{t}$ represents a stage of the learning process that can be warm-started given the outcome of the previous stage. This is particularly straightforward for neural network training: given a space of weights $\mathcal{W}$, each $A_{t}:\mathcal{W}\times\Lambda\to\mathcal{W}$ could be, for instance, an epoch of gradient-based optimization of the network weights (given configuration hyperparameters). Let $N\in\mathbb{N}$ be a population size and $\{(\theta_{0}^{j},\lambda_{0}^{j})\}_{j\in[N]}$ be a set of initial model weights and hyperparameters, where $\theta_{0}^{j}\in\mathcal{W}$ and $\lambda_{0}^{j}\in\Lambda$. These pairs represent the “individuals” in PBT and they can be initialized drawing from suitable distributions.
In PBT, generations correspond to stages of the learning algorithm. At each generation $t\in[T]$ the weights are updated as follows
$\tilde{\theta}_{t}^{j}=A_{t}(\theta_{t-1}^{j},\lambda_{t-1}^{j})\quad\text{for}j\in[N].$
Then, evolutionary transformations are applied to both parameters and hyperparameters before proceeding to the subsequent learning stage, on the basis of a fitness function $\mathscr{M}$ (e.g. a validation error) evaluated on the current model weights. Suppose we sort individuals in ascending order so that $\mathscr{M}(\tilde{\theta}_{t}^{1})\leq\mathscr{M}(\tilde{\theta}_{t}^{2})\leq\dots\leq\mathscr{M}(\tilde{\theta}_{t}^{N})$. jaderberg2017population propose two type of operations:
@@
  • exploit, a selection strategy that involves copying the parameters and hyperparameters of better performing individuals onto worst performing ones; for instance, one may set $(\theta^{N}_{t},\lambda^{N}_{t})=(\tilde{\theta}^{1}_{t},\lambda^{1}_{t-1})$. jaderberg2017population suggest to replaces the worst-performing 20% with individuals randomly drawn from the best-performing 20%;
  • explore, a mutation transformation that brings over the parameter form the last learning stage, but either resamples the hyperparamter from the starting distribution, or perturbs them, e.g. setting $(\theta_{t}^{j},\lambda_{t}^{j})=(\tilde{\theta}_{t}^{j},\lambda_{t-1}^{j}+\varepsilon)$ for a suitable perturbation $\varepsilon$.
PBT further uses elitism, leaving the top performing individuals unchanged; that is, setting for some $M\in[N]$, $(\theta_{t}^{i},\lambda_{t}^{i})=(\tilde{\theta}_{t}^{i},\lambda_{t-1}^{i})$ for $i\in[M]$. Figure 6.1 present a visualization of a phylogenetic tree of game-playing policy functions for Atari games. The performance metric is the expected episodic reward, and the models deep neural networks trained using the Feudal Networks (FuN) framework [vezhnevets2017feudal].
@@
tao2020learning combine PBT with gradient-based optimization of the hyperparameters, which we will discuss in the next chapter.